# A Survey on Text-guided 3D Visual Grounding: Elements, Recent Advances, and Future Directions

### v1

@@ 4.5 Discussion
5 Performance Comparison
We have summarized the performance of various already-published T-3DVG methods on the ScanRefer, Sr3D, and Nr3D benchmark datasets since 2021 in Tables III, IV, and V in chronological order, respectively. In these tables, we also list the framework type (two-stage or one-stage), the input modality, and the supervision method (fully-supervised or weakly-supervised) for each method.
Overall, two-stage methods tend to perform slightly worse than one-stage methods in grounding due to their high dependency on the performance of pretrained 3D detection models, which can overlook unimportant or inconspicuous objects. However, some recent two-stage methods [86, 22] have addressed this shortcoming by refining the detected object locations through text guidance or other strategies. For the same method, results are always better when the input includes additional multiview 2D information compared to those using only the 3D modality as input. In the tables, it can also be observed that methods employing multi-task joint training [55, 87, 56] often achieve relatively better results. This improvement can be attributed to the benefits of more data and the learning of more comprehensive feature representations brought by learning from different tasks.
By comparing the results across three datasets, it can be observed that the relative performance of methods is generally consistent across different datasets. However, there are subtle differences between the tasks of the ScanRefer dataset and the two datasets of Referit3D (Sr3D, Nr3D). ScanRefer dataset requires methods to regress the bounding box of the target object, while Referit3D datasets only require selecting among the provided bounding boxes. Therefore, some one-stage methods have a more significant advantage on ScanRefer due to their ability to achieve higher quality bounding boxes, whereas this advantage is less noticeable on Sr3D and Nr3D dataset. The comparison between EDA [26] and ViL3DRel [22] provides a clear example of this. Note that the results of the CORE-3DVG [86] on the Referit3D dataset are based on the detected bounding boxes for grounding, rather than the conventional ground truth bounding boxes, resulting in slightly inferior performance compared to the results on the ScanRefer dataset.
Additionally, to ensure a fair comparison, we have listed all methods involving large language models (LLMs) in another Table VI, regardless of whether LLMs are used for data pre-processing or direct training of multi-modal LLMs. As can be seen from the table, due to the varying use of LLMs, the results of these methods differ significantly. Methods that employ instruction-tuning with multimodal LLMs [100] perform worse compared to common T-3DVG methods; conversely, using LLMs for commonsense reasoning on processed information [113] yields impressive results. This variance in performance underscores the complexity and potential of integrating LLMs in 3D visual grounding tasks.
6 Challenges and Future Directions
6.1 Critical Analysis
Relying on a large amount of annotations. Existing methods for the T-3DVG task can be generally categorized into two types based on task settings: fully-supervised [7, 8, 50, 65, 29, 27, 55, 26, 83, 69] and weakly-supervised methods [31, 106]. However, under the fully-supervised setting, a huge amount of bounding boxes for numerous objects in the 3D point cloud scenes, along with their descriptive texts, are required to be annotated as labels to provide reliable supervision. For example, the widely used ScanRefer dataset [7] contains 806 scenes with 51,583 matched object-text labels. However, manually annotating these object-text pairs is very time-consuming and labor-intensive. To alleviate this problem, the weakly-supervised methods [31, 106] are proposed to only use text labels for grounding without relying on any object bounding box annotations. In the weakly supervised setting, we do not need to annotate bounding boxes for the objects, but descriptive texts still need to be annotated for each object in every scene. Although they slightly reduce the annotation cost, tens of thousands of language descriptions still need to be annotated, requiring significant manual effort. Therefore, how to learn a 3D visual grounding model with as few annotations as possible is essential to investigate.
@@ 6.2 Future Directions
Group-wise 3D grounding. Despite the significant advancements made in T-3DVG field, the existing 3D grounding methods are overly idealistic, which aim to localize a unique target object that has been known to exist in a single 3D scene. This setting limits the applicability of 3D object grounding in real-world scenarios. In general, a group of related 3D scenes is available (e.g., a house comprises several rooms), some of which may not contain the described object while some may potentially contain multiple target objects. Therefore, it is essential to develop and address a new realistic setting, namely Group-wise 3D Object Grounding, where a flexible number of target objects should be localized in a group of related 3D scenes given a referring sentence. The group-wise 3D object grounding setting should generalize the traditional 3D object grounding task in the following two aspects: (1) Compared with the traditional single target assumption, the new setting allows for the existence of zero or multiple target objects in a specific 3D scene. (2) The new setting requires querying a group of related 3D scenes simultaneously with a single referring expression. Note that the application scope of group-wise 3D object grounding is the superset of the traditional task, which means (1) or (2) can be omitted based on the actual situations, and thus our setting is more natural and practical in broader real-world scenarios. Potential solutions [132, 133, 134, 135] can devise a two-stage framework that separately locates objects in each scene or design a one-stage framework that jointly retrieval all objects.
@@
Better integration with large models. Multi-modal Large Language Models (MLLMs) have achieved significant success and demonstrated promising capabilities in various 2D multimodal downstream tasks, such as text-to-image generation [136, 137, 138], visual question-answering [139, 140, 141], and etc., due to an increase in the amount of data, computational resources, and number of model parameters. By further benefiting from the strong comprehension of large language models (LLMs) [142, 143, 144], recent 2D MLLMs [145, 122, 123] on top of LLMs show superior performances in solving complex vision-language tasks by utilizing appropriate human-instructed prompts. By applying the LLMs and pre-trained large 3D vision-language models into the T-3DVG field, it can also provide more representative multimodal features and more accurate cross-modal alignment for complicated scene-based understanding and reasoning. There are many recently proposed 3D large language models [146, 147, 81, 148], showing a strong power in handling 3D reasoning tasks. However, their proposed methods simply follow a general and unified framework that can only address the global tasks, such as classification, captioning, generation, and question answering. Their performances are poor on the T-3DVG as this task requires detailed local reasoning. Although some existing T-3DVG works [44, 112, 113, 96] have tried to employ LLMs to individually comprehend and enrich the textual semantics, they still ignore investigating the strong ability of 3D large models to handle multimodal inputs. Therefore, how to incorporate large models into the specific T-3DVG grounding task with in-depth structure level designs is an emergency issue.
7 Conclusion
@@
Text-guided 3D visual grounding (T-3DVG) is a fundamental but challenging vision-language task in multimedia understanding. It is also worth exploring since it can serve as an intermediate step for various downstream tasks, such as multi-modal 3D scene understanding, language-guided robotic grasping, and 3D embodied interaction. In this survey, we comprehensively review the fundamental elements and recent research advances of T-3DVG. To be more specific, we first provide a general structure of the T-3DVG pipeline with detailed components in a tutorial style. We then summarize the existing T-3DVG approaches into different categories and analyze their strengths and weaknesses. After that, we present the benchmark datasets and evaluation metrics to assess their performances. Finally, we discuss the potential limitations of existing T-3DVG and figure out some insights on future research. We hope this survey can provide handy guidance when designing T-3DVG models and attract more researchers to contribute to this field.

### v2

@@ V Performance Comparison
VI Challenges and Future Directions
Despite significant performances achieved on benchmark datasets, limitations of T-3DVG methods hinder the practicability and generalization ability of T-3DVG methods. In this section, we first provide the critical analysis of current methods and then discuss potential future directions.
VI-A Critical Analysis
Relying on a large amount of annotations. Most existing methods for the T-3DVG task are implemented under a fully-supervised setting. However, in this setting, a huge amount of bounding boxes for numerous objects in the 3D point cloud scenes, along with their descriptive texts, are required to be annotated as labels to provide reliable supervision. For example, the widely used ScanRefer dataset [11] contains 806 scenes with 51,583 matched object-text labels. However, manually annotating these object-text pairs is very time-consuming and labor-intensive. To alleviate this problem, the weakly-supervised methods are proposed to only use text labels for grounding without relying on any object bounding box annotations. In the weakly-supervised setting, we do not need to annotate bounding boxes for the objects, but descriptive texts still need to be annotated for each object in every scene. Although the weakly-supervised methods slightly reduce the annotation cost, tens of thousands of language descriptions still need to be annotated, requiring significant manual effort. Therefore, how to learn a 3D visual grounding model with as few annotations as possible is essential to investigate.
@@ VI-B Future Directions
Better integration with large models. Multi-modal Large Language Models (MLLMs) have achieved significant success and demonstrated promising capabilities in various 2D multimodal downstream tasks, such as text-to-image generation [150, 151, 152], visual question-answering [153, 154, 155], and etc., due to an increase in the amount of data, computational resources, and number of model parameters. By further benefiting from the strong comprehension of large language models (LLMs) [156, 157, 158], recent 2D MLLMs [159, 128, 129] on top of LLMs show superior performances in solving complex vision-language tasks by utilizing appropriate human-instructed prompts. By applying the LLMs and pre-trained large 3D vision-language models into the T-3DVG field, it can also provide more representative multimodal features and more accurate cross-modal alignment for complicated scene-based understanding and reasoning. There are many recently proposed 3D large language models [160, 161, 142, 162], showing a strong power in handling 3D reasoning tasks. However, their proposed methods simply follow a general and unified framework that can only address global tasks, such as classification, captioning, generation, and question answering. Their performances are poor on the specific T-3DVG as this task requires detailed local reasoning. Although some existing T-3DVG works [50, 116, 117, 97] have tried to employ LLMs to individually comprehend and enrich the textual semantics, they still ignore investigating the strong ability of 3D large models to handle multimodal inputs. Therefore, how to incorporate large models into the specific T-3DVG grounding task with in-depth structure-level designs is an emergency issue.
VII Conclusion
@@
Text-guided 3D visual grounding (T-3DVG) is a fundamental yet challenging vision-language task in multimedia understanding. It is also worth exploring since it serves as an intermediate step for various downstream tasks, such as multi-modal 3D scene understanding, language-guided robotic grasping, and 3D embodied interaction. In this survey, we comprehensively review the fundamental elements and recent research advances of T-3DVG. Specifically, we first provide a general structure of the T-3DVG pipeline with detailed components in a tutorial style. We then summarize the existing T-3DVG approaches into different categories and analyze their strengths and weaknesses. Afterwards, we present the benchmark datasets and evaluation metrics to assess their performances. Finally, we discuss the potential limitations of existing T-3DVG and provide some insights in future research.